OpenStudy (vera_ewing):

What is an equation of the line, in point-slope form, that passes through the given points and has the given slope. Point: (8, -8) Slope: 3

OpenStudy (vera_ewing):

@phi i really need help with this

OpenStudy (cj49):

y-y1=3(x-x1) y-(-8)=3(x-8)

OpenStudy (vera_ewing):

thanks so much! can you help with one more? :)

OpenStudy (cj49):

ull need to simplify that equation...sure